The sum of two numbers is -7. Four times first number plus twice the second number is -8. Find the two numbers.
Answer by rfer(16322)

x+y=-7
x=-y-7
---------
4x+2y=-8
4(-y-7)+2y=-8
-4y-28+2y=-8
-4y+2y=28-8
-2y=20
y=20/-2
y=-10
-----------------
4x+2(-10)=-8
4x-20=-8
4x=-8+20
x=12/4
x=3
----------------------
4*3+2*-10=-8
12-20=-8